In a plate glass factory, sheets of glass move along a conveyor belt at a speed of 15.0 cm/s. An automatic cutting tool descends at present intervals to cut the glass to size. since the assembly belt must keep moving at constant speed, the cutter is set to cut at an angle to compensate for the motion of the glass. If the glass is 72.0 cm wide and the cutter move across the the width at a speed of 24.0 cm/s , at what angle should the cutter be set . according the book the answer is 32.0 grades but I don't know how to get that answer.
Answer #1

time taken by cutter to cut , t = w/vcutter = 72/24 = 3 s


distance moved by glas in 3s , x = 15*3 = 45 cm

tantheta = x/w = 45/72

theta = tan^-1(45/72) = 32 degrees <<-----------ANSWER